### v1

# Convolutional autoencoders for the reconstruction of three-dimensional interfacial multiphase flows

Abstract
In this work, we perform a comprehensive investigation of autoencoders for reduced-order modeling of three-dimensional multiphase flows. Focusing on the accuracy of reconstructing multiphase flow volume/mass fractions with a standard convolutional architecture, we examine the advantages and disadvantages of different interface representation choices (diffuse, sharp, level set). We use a combination of synthetic data with non-trivial interface topologies and high-resolution simulation data of multiphase homogeneous isotropic turbulence for training and validation. This study clarifies the best practices for reducing the dimensionality of multiphase flows via autoencoders. Consequently, this paves the path for uncoupling the training of autoencoders for accurate reconstruction and the training of temporal or input/output models such as neural operators (e.g., FNOs, DeepONets) and neural ODEs on the lower-dimensional latent space given by the autoencoders. As such, the implications of this study are significant and of interest to the multiphase flow community and beyond.
keywords:
autoencoder , multiphase flow , reduced order model , deep learning , interface capturing
[…]
[label1]organization=Department of Mechanical Engineering, Stanford University, city=Stanford, country=USA
[…]
[label2]organization=Department of Engineering Mechanics, KTH Royal Institute of Technology, city=Stockholm, country=Sweden
1 Introduction
Fluid flows are considered high-dimensional dynamical systems, where real-time prediction, control, and optimization are prohibitively expensive without some form of reduced-order modeling. A low-dimensional representation of the state of a fluid system is required for a reduced-order model (ROM). While there has been substantial progress in developing ROMs for single-phase flows [1, 2, 3, 4], interfacial multiphase flows present a unique challenge. Multiphase flows consist of two or more immiscible fluid phases separated by a deformable interface. In this case, the state of the dynamical system strongly depends on the instantaneous location of the different phases, due to the discontinuity in density, viscosity, pressure, and other fields across the deformable phase interface. As such, it is crucial for the ROM to represent the shape of three-dimensional interfaces in multiphase flows accurately. In this work, we study the performance of convolutional autoencoders (AE) for this task. We reveal the impact of the interface representation type (diffuse, sharp, level-set function), shape complexity, training set set size, and compression ratio on the performance of the AE. This work constitutes a robust building block for developing ROMs to predict temporal dynamics in multiphase flows. Finally, considering that low-dimensional representations of complex three-dimensional shapes are highly sought after in many applications, including computer graphics, computer vision, and robotics, our findings may have implications beyond multiphase flow modeling [5, 6, 7].
While substantial progress has been made in simulating (full-order modeling) interfacial multiphase flows using various approaches in the past few decades [8, 9], these simulations are too expensive to allow for real-time predictions, control, or optimization. As such, ROMs for such multiphase flows are sought after. Such models are scarce in the literature [10, 11, 12, 13, 14], especially when one seeks ROMs that aim to capture the interfaces and their evolution in such flows. Here, we focus on the low-dimensional representation of the interfaces in multiphase flows via convolutional autoencoders (AEs). An AE is a neural network consisting of an encoder that maps the input onto a latent representation and a decoder that reconstructs the input from this representation. The encoder performs a nonlinear transformation, often reducing dimensionality, while the decoder attempts to approximate the original input as closely as possible. AEs have been successful in low-dimensional representation of single-phase flows [15, 16, 17, 18], but their application to the reconstruction of multiphase flow states remains limited [10, 13]. There are multiple ways to implicitly represent interfaces in multiphase flows via field variables. These include diffuse interface representations like phase field variables, sharp interface representations like volume fractions in Volume-of-Fluid models, and level-set functions (signed-distance functions), which are inherently available in level-set-based multiphase flow models. Since one can transform these representations into one another (e.g., as a pre-processing step), we explore how the interface representation choice influences the performance of the AE. For training, we use a series of synthetic data sets with increasing interface complexity, as well as a data set consisting of snapshots from high-fidelity simulations of drops in decaying homogeneous isotropic turbulence (HIT) [14]. This also allows us to study the effect of the complexity of the interfacial structures on the reconstruction accuracy of the AEs. All in all, accurately representing flow states in a low-dimensional space is a prerequisite for most ROMs. This work elucidates the best practices and limitations of applying AEs for this purpose in multiphase flows.
Figure 1: Illustration of the three interface representations compared in this study: (a) signed-distance (level-set) functions, (b) a diffuse (tanh) interface profile, and (c) a sharp-interface indicator function, showing the interface in one snapshot of the simulation of a droplet in homogeneous isotropic turbulence.
2 Methodology
2.1 Problem formulation
We seek to train convolutional autoencoders (AEs) for the reconstruction of interfacial structures in two-phase systems. Interfacial structures can be captured with various field variables. Figure 1 shows three different choices for full-order representation of the same interfacial structure, where panel (a) displays the signed distance function (SDF), also known as the level-set function, typically denoted with $s$. Panel (b) displays a diffuse interface representation using a so-called phase field variable given by
$\phi=\frac{1+\tanh\left(\frac{s}{2\epsilon}\right)}{2},$ (1)
where $\epsilon$ is the chosen interface thickness. We denote a diffuse interface representation with thickness $\epsilon$ with “Tanh $\epsilon$”. Panel (c) gives a sharp interface representation, with an indicator function defined as
$H=\frac{1-\operatorname{sgn}(s)}{2}.$ (2)
While these different fields represent the same interface (or phase) distributions, it is clear from Figure 1 that they have vastly different properties. Let us denote these fields, $(s,H,\phi)$, collectively as a three-dimensional input field $x$. The autoencoder, composed of an encoder $f_{\theta_{e}}^{e}$ and a decoder $f_{\theta_{d}}^{d}$, where $\theta_{e}$ and $\theta_{d}$ denote the learnable parameters of the encoder and decoder networks, respectively, is trained to reconstruct $x$ as accurately as possible by minimizing a loss function that quantifies the reconstruction error. Two common choices for this loss function are the mean squared error (MSE):
$\mathcal{L}(x,\hat{x})=\|x-\hat{x}\|_{2}^{2},$
and the mean absolute error (L1):
$\mathcal{L}(x,\hat{x})=\|x-\hat{x}\|_{1},$
where $\hat{x}=f_{\theta_{d}}^{d}(f_{\theta_{e}}^{e}(x))$ is the reconstructed output. The training process therefore amounts to solving the optimization problem
$\min_{\theta_{e},\theta_{d}}\mathcal{L}\left(x,f_{\theta_{d}}^{d}\left(f_{\theta_{e}}^{e}(x)\right)\right),$ (3)
where $\mathcal{L}(\cdot,\cdot)$ denotes the chosen loss function. Note that $f_{\theta_{e}}^{e}(x)=z$ is the latent representation of the input, and the compression ratio is defined as the ratio of the dimensionality of the input (or output) field to that of the latent variable.
2.2 Dataset
(a) HIT simulation
(b) Synthetic ($\mu=1$)
(c) Synthetic ($\mu=2$)
(d) Synthetic ($\mu=2.5$)
Figure 2: Visualization of the interfacial contour for 3 samples from each of the four datasets considered in this work. In (a) the HIT simulation dataset is shown, while the remaining three panels show the synthetic dataset which is parameterized by $\mu$. $\mu=1$ in (b), $\mu=2$ in (c), and $\mu=2.5$ in (d). $\mu$ parametrizes these datasets through Equation 5.
Two datasets are used in this work. Each dataset is divided into a training split, a test split, and a hyper-parameter validation split in a 80/15/5 ratio.
2.2.1 Interfacial flow simulation dataset
To incorporate a realistic flow, we use snapshots from simulations of drops in decaying homogeneous isotropic turbulence (HIT) that we previously used for training a machine learning (ML) model for drop breakup prediction [14]. The simulations were performed on a $256^{3}$ grid using a phase field (diffuse interface) approach [19, 20]. Patches of size $64^{3}$ were extracted from these snapshots, by randomly choosing 64 patch locations per volume and discarding empty patches. Knowing the interface thickness ($\epsilon_{\text{sim}}$) used in the simulations, we compute the SDF ($s$ in Section 2), from which we construct input fields with various diffuse interface thicknesses, given by
$\phi=\frac{1+\tanh\left(\frac{s}{2\epsilon}\right)}{2},$ (4)
with the limit of $\epsilon\to 0$ giving us the sharp interface representation.
2.2.2 Synthetic dataset
To study the effect of interface complexity, we construct a series of three synthetic datasets, which are illustrated in Figure 2. We parameterize the data generation using a single variable, denoted by $\mu$, which controls the droplet size distribution and number of droplets per sample. In this way, as $\mu$ decreases, we obtain synthetic datasets with progressively increasing topological and geometrical complexity, and finer spatial scales. Each sample is constructed by taking the union of $N_{D}$ spherical droplets, where the droplet radii $r$ are drawn independently from
$r\sim\mathrm{lognormal}(\mu,\,1/2).$ (5)
$N_{D}$ is then drawn from a uniform distribution such that the total volume fraction is uniformly distributed in $[0,1]$. All samples are $64^{3}$.
2.3 Neural network architecture
[…]
Figure 3: Default residual block used in the autoencoder.
Throughout this work, we adopt a standard convolutional architecture; our objective is to compare interface representations rather than architecture optimization. The choice of a fully convolutional architecture [21] is motivated by our application to (potentially large) 3D volumes, as it enables a patch-based approach due to the translation equivariance of convolutional layers. Our network architecture follows the standard encoder-decoder structure described in Equation 1 and is based on the residual block illustrated in Figure 3. Each layer in the encoder (decoder) consists of two such blocks in series, followed by a downsampling (upsampling) operation. The choice of a ResNet-type architecture [22] along with weight-standardized convolution layers is advantageous in the small batch size regime [23], which is necessarily the case in 3D due to the large feature maps which must be stored.
Given an input of shape $1\times H\times W\times D$, we obtain a latent representation of shape $Z\times\frac{H}{2^{N}}\times\frac{W}{2^{N}}\times\frac{D}{2^{N}}$, where $N$ is the number of downsampling layers in the encoder. Unless otherwise specified, we use $N=4$ and $Z=4$, resulting in a compression ratio of $1024$ and approximately $5.4\times 10^{6}$ trainable parameters.
2.4 Training
All models were trained using the Adam optimization algorithm [24] and mean absolute error loss, with a batch size of 4 and a learning rate of $10^{-5}$ (see hyper-parameter study in Section 3.3). The number of training epochs was held constant at 100 for the synthetic datasets and 15 for the HIT simulation dataset. This was confirmed qualitatively to provide sufficient convergence. Each training run utilised 4 Nvidia V100 (16GB) GPUs in a data-parallel approach.
3 Results and discussion
3.1 Evaluation metrics
Each case is evaluated using a held-out test set, according to the splits described in Section 2.2. Rather than standard $L^{p}$ error norms, we adopt two metrics from the field of image segmentation. The evaluation of a (semantic) image segmentation is highly similar to the evaluation of an interfacial geometry. In both cases, the prediction consists of an arbitrary region represented by a binary mask. We transform all predictions to a sharp interface representation before computing metrics. The following two metrics offer a concise and understandable summary of the performance of this prediction.
First, the Dice coefficient [25] provides a value in $[0,1]$ describing the overall volumetric agreement, where a value of 1 is the best possible score:
$\mathrm{DSC}(X,Y)=\frac{2\,|X\cap Y|}{|X|+|Y|}.$ (6)
Second, the Hausdorff distance [26] is a worst-case measure, analogous to the $L^{\infty}$ norm. It is the maximum distance between the predicted and the true interface:
$\mathrm{HD}(\Gamma_{X},\Gamma_{Y})=\max\left\{\sup_{x\in\Gamma_{X}}\inf_{y\in\Gamma_{Y}}|x-y|,\;\sup_{y\in\Gamma_{Y}}\inf_{x\in\Gamma_{X}}|x-y|\right\},$ (7)
where $\Gamma_{X}$ represents an interface. We normalize this distance relative to the length scale of the domain, so a value of 1 corresponds to a distance of 64 cell widths (for a $64^{3}$ grid), and a value of 0 corresponds to the best possible score.
3.2 Training process uncertainty
Due to computational cost constraints (each full training run requires approximately 40 GPU-hours) we are unable to repeat each result with multiple seeds (for the random weight initialization and batch order during training). Instead, we measure the effect of the random seed once, for three different interface representations and five different random seeds, using 50% of the HIT training set for efficiency purposes. The uncertainty in the training process is summarized in Table 1. Overall, these uncertainties are smaller than many of the differences between interface representation methods measured in subsequent results.
Dice Hausdorff distance
Interface Mean Std 95% CI Mean Std 95% CI
SDF 0.867 0.025 (0.836, 0.899) 0.185 0.127 (0.0274, 0.342)
Sharp 0.953 0.004 (0.948, 0.958) 0.0577 0.0090 (0.0465, 0.0689)
Tanh 1/32 0.957 0.004 (0.951, 0.962) 0.0666 0.0069 (0.0580, 0.0752)
Table 1: Measurement of training uncertainty. Mean, standard deviation and 95% confidence interval (CI) for Dice and Hausdorff distance, based on repeats with five random seeds. 50% of HIT simulation dataset used.
3.3 Hyper-parameter study
A limited hyper-parameter grid search was run using the HIT simulation dataset. The search space consisted of learning rate, loss function, weight decay, and activation function and is summarized in Table 2. The dice scores on the hyper-parameter validation split for three interface types (sharp, tanh $\frac{1}{32}$, SDF) are summarized by a parallel coordinate plot in Figure 4.
Hyper-parameter Grid search values
Loss Function $\bm{L_{1}}$, MSE
Learning Rate $10^{-3}$, $10^{-4}$, $\bm{10^{-5}}$
Weight Decay ($L_{2}$) $10^{-4}$, $\bm{10^{-6}}$
Activation Function SiLU, ReLU, Tanh
Table 2: Hyper-parameter grid search space. Optimal values used in the remainder of this work highlighted in bold.
The optimal weight decay and activation functions are unambiguously $10^{-6}$ and SiLU. The optimal learning rate is $10^{-5}$, but the sharp interface representation is also tolerant of a $10^{-4}$ learning rate. The choice of loss function appears to be interface-type dependent, as shown in Figure 5. For sharp representations, both MSE and $L_{1}$ are within the uncertainties estimated in the previous section. However for the more diffuse tanh interface representations and the SDF, using the $L_{1}$ loss offers a distinct advantage. This result may be understood in terms of MSE loss providing a smaller signal close to the interface for smoother representations. Further details on the hyperparameter search are provided in A.
(a) Sharp interface
(b) Tanh $\frac{1}{32}$ interface
(c) SDF interface
Figure 4: Parallel coordinate plot showing hyper-parameter grid search results for three different interface representations.
Figure 5: Hyper-parameter grid search results examining the effect of loss function for two learning rates. The $10^{-3}$ learning rate results are omitted as they did not converge. The activation function and weight decay are held at their optimal values here (SiLU and $10^{-6}$ respectively).
3.4 Performance vs. interface representation
(a) HIT simulation dataset
(b) Synthetic datasets
Figure 6: Absolute performance of each interface representation on (a) HIT simulation dataset and (b) synthetic datasets. Recall that for Dice, higher is better, while for Hausdorff, lower is better.
(a) HIT simulation dataset
(b) Synthetic datasets
Figure 7: Relative performance of each interface representation on (a) HIT simulation dataset and (b) synthetic datasets. Recall that for Dice, higher is better, while for Hausdorff, lower is better.
(a) Tanh $\frac{1}{32}$
(b) Tanh $\frac{1}{128}$
Figure 8: Visualization of a single test set case from the synthetic ($\mu=1$) dataset with (a) smoother interface and (b) sharper interface.
Figures 6 and 7 show the absolute and relative reconstruction performance for each dataset using various interface representations. The Dice coefficient is generally lowest for the SDF inputs and steadily improves as the interface becomes sharper, plateauing at a dataset-dependent interface thickness $\epsilon$.
While the Dice metric favors sharper interfaces, a different trend emerges when evaluating the Hausdorff distance. For more complex fields containing small drops, the sharpest representations no longer yield the best performance. Instead, a moderately diffuse interface representation with $\epsilon\approx\frac{1}{32}$ tends to give the lowest Hausdorff error. A representative example shown in Figure 8 illustrates this effect: sharper interfaces enable more accurate reconstruction of large-scale features, but at the expense of small-scale ones. We attribute this trade-off to the spectral bias of deep neural networks [27], which makes high-frequency features (such as sharp interfaces and fine structures) more difficult to learn during training.
The droplet size distribution is an important quantity of interest in many interfacial multiphase flows. In Figure 9, we compare the predicted distribution of droplet size in the reconstructed samples on the synthetic dataset. In particular, when we focus on the most complex dataset ($\mu=1$), we observe that the best match to the ground truth was given by a tanh profile with an intermediate interface thickness. This observation further supports the conclusion that moderately diffuse interfaces offer the best balance between resolving large and small features.
The underperformance of SDFs was unexpected and highlights an important distinction between interface representations. We hypothesize that this is because reconstruction errors in the SDF representation are distributed throughout the domain, rather than being localized near the interface, and thus do not penalize the model as effectively as errors in sharp or diffuse representations. Overall, our findings suggest a trade-off between sharp and diffuse representations: while sharper interfaces improve the reconstruction of large-scale features, moderately diffuse representations better capture small-scale structures in complex fields. Despite this trade-off, both sharp and diffuse representations consistently outperform SDFs. We repeated these experiments at compression ratios ranging from 100 to 8000 (not shown for brevity), and observed that these trends persist across compression levels.
Figure 9: Comparison of predicted vs. true droplet size distributions for different interface representations. On the most complex dataset ($\mu=1$), only the tanh representation with intermediate interface thicknesses provides predictions which preserve the droplet size distribution.
3.5 Performance vs. dataset size
For each interface representation and dataset, we investigate the effect of progressively reducing the training set size. As shown in Figure 10, the model performance on less complex datasets continues to improve with more data, converging toward perfect reconstruction at a rate of $\mathcal{O}(N^{-\frac{1}{2}})$. In contrast, for the most complex datasets ($\mu=1$ and $\mu=2$), the model architecture, rather than the dataset size, appears to be the limiting factor.
(a) HIT simulation dataset
(b) Synthetic datasets
Figure 10: Performance versus training set size for each dataset.
3.6 Generalization error
Finally, we investigate the generalization ability of each interface representation by evaluating the mean Dice coefficient for each combination of train/test dataset. Figure 11 shows that sharper interface representations transfer better to out-of-domain test datasets. Furthermore, sharp interface models trained only on synthetic data transfer surprisingly effectively to the HIT dataset, suggesting that synthetic data augmentation is an effective tool in this domain.
Figure 11: Comparison of generalization error for different interface representations when training and evaluating on different combinations of datasets.
4 Conclusions
We have evaluated the performance of a fully convolutional 3D autoencoder in reconstructing interfacial geometries using common implicit interface representations employed in multiphase simulations. Our comparisons across simulated and synthetic datasets of varying complexity revealed that the signed distance function consistently yields poorer reconstruction performance. In contrast, sharp and diffuse interface representations exhibit a trade-off: sharper interfaces improve the reconstruction of large-scale features, but excessively sharp representations tend to miss small-scale structures. On the other hand, overly diffuse interfaces lead to degraded overall accuracy.
Overall, we find that a moderately diffuse interface representation, with an interface width of 141-4 times the grid spacing, offers the best balance across datasets and metrics, preserving small features while maintaining high reconstruction fidelity.
Future work will focus on improving the autoencoder by incorporating conservation constraints on the network output, exploring variational autoencoders, and including additional state variables relevant to multiphase flows. Moreover, the insights gained from this comprehensive analysis lay the groundwork for developing models capable of predicting temporal dynamics in multiphase systems.
Acknowledgements
We acknowledge financial support from the US Department of Energy’s National Nuclear Security Administration via the Stanford PSAAP-III Center (DE-NA0003968). S. M. also acknowledges start-up funding from KTH Royal Institute of Technology and the Swedish e-science Research Center.
Appendix A Best performing hyper-parameter sets
We provide further detail on the optimal hyper-parameters for each interface type here. Figure 12 lists the top 5 hyper-parameter sets examined in the grid search experiment for four of the interface types.
(a) Tanh $\frac{1}{8}$
(b) Tanh $\frac{1}{128}$
(c) Sharp
(d) SDF
Figure 12: Top 5 best performing hyper-parameter sets for four interface types.